OpenStudy (anonymous):

Use the definition of the derivative to find f prime(x) for: f(x)= x^2-2x

OpenStudy (anonymous):

Please help!

OpenStudy (anonymous):

2x-2

OpenStudy (anonymous):

use nx^n-1 but yeah akamushi is right

OpenStudy (anonymous):

if you're going to use the definition you'll need to find the limit of the difference quotient: \[\large \lim_{h \rightarrow 0} \frac {f(x+h) - f(x)}{h}\] \[\large \lim_{h \rightarrow 0} \frac {[(x+h)^2-2(x+h)] - [x^2-2x]}{h}\] \[\large \lim_{h \rightarrow 0} \frac {x^2+2xh+h^2-2x-2h - x^2+2x}{h}\] \[\large \lim_{h \rightarrow 0} \frac {2xh+h^2-2h}{h}\] \[\large \lim_{h \rightarrow 0} \frac {h(2x+h-2)}{h}\] \[\large \lim_{h \rightarrow0} 2x+h-2\] \[\large =2x+0-2 = 2x-2\]

OpenStudy (anonymous):

Thank you, this help me a lot more.